OpenStudy (anonymous):

Lol I forgot how t divide polynomials by binomials. How do you solve one like x^3+6x+3 all over x+3.

OpenStudy (anonymous):

Typos. *how to *x+3?

OpenStudy (anonymous):

x+3=0 x=-3 f(x)=x^3+6x+3 (-3)^2+6(-3)+3=-27-18+3=-42

OpenStudy (anonymous):

So, thats how you divide it? By getting the f(x)?

OpenStudy (anonymous):

Long Division dude!

OpenStudy (anonymous):

It's hard to write it here, but this link explains it well : http://www.purplemath.com/modules/polydiv2.htm